OpenStudy (anonymous):

Which numbers are greater than 7x10^-4? Choose all answers that are correct. A. 8x 10 ^-4 B. 9x10^-5 C. 4x10^-4 D. 3x10^-2 E. 7x10^-5

OpenStudy (jhannybean):

You've got \[7 \cdot 10^{-4}\] When a number is raised to the -ve power, it can be written as a fraction. say x is any real number, therefore \(x^{-1} \implies \frac{1}{x}\) Keeping that in mind, we can rewrite your function as \[7 \cdot \frac{1}{10^4}\]Now all you have to do is expand 10^4 and divide it by 7 to get your answer.

OpenStudy (anonymous):

Just give me a simple answer plz lol.

OpenStudy (jhannybean):

Let's work it out together. What is 10^4?

OpenStudy (anonymous):

I'm a politician, not a mathlete.

OpenStudy (jhannybean):

\[ 10^4 = 10\cdot10\cdot10\cdot10\] Count up all the 0's and stick a 1 infront, what do you get?

OpenStudy (anonymous):

Refer to the message above.

OpenStudy (anonymous):

This site is broken. Lol